Find the zeros:
f(x)= sin2xsinx-cosx

Thanks

Answer by lwsshak3(11628)

Find the zeros: For [0,2π)
f(x)= sin2xsinx-cosx =2sinxcosxsinx-cosx=cosx(2sin^2x-1)=0
cosx=0
x=π/2, 3π/2
or
2sin^2(x)-1=0
sin^2(x)=1/2
sinx=±√(1/2)=±√2/2
x=π/4,3π/4,5π/4,7π/4